ok the question is 1/5x+1/4=3 and the correct answer is 65/4 can u please help me step by step and show me how this came to be the correct answer
Found 2 solutions by richwmiller, nikolatesla1:
Answer by richwmiller(17219)

1/5x+1/4=3
could be
(1/5)*x+1/4=3
x/5=2 3/4
x/5=11/4
x=55/4

then x=55/4 not 65/4
or
1/(5x)+1/4=3
then x=4/55

Answer by nikolatesla1(6)

1/5x+1/4=3. subtract 1/4 from both sides to get 1/5x=2.75; then divide both sides by 1/5 to get x by itself to get x=55/4